# CoSQA+: Enhancing Code Search Dataset with Matching Code

### v1

@@ Abstract
[…]
tcb@breakable
CoSQA+: Enhancing Code Search Dataset with Matching Code
Jing Gong1thanks:  Both authors contributed equally to this work., Yanghui Wu [^1] footnotemark: 1, Linxi Liang1, Zibin Zheng2, Yanlin Wang3thanks:  Corresponding author. 1, 2, 3Sun Yat-sen University, Guangzhou, Guangdong, China 1{gonj39,wuyh223,lianglx26}@mail2.sysu.edu.cn 2zhzibin@mail.sysu.edu.cn, 3yanlin-wang@outlook.com
1 Introduction
@@
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al., 2013; Sim et al., 2011; Stolee et al., 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al., 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. For the practical usability of neural code search, real user queries are crucial to a high-quality dataset  (Sun et al., 2022). The key challenge is how to pair real-world queries with codes for a high-quality code search benchmark.
However, there are some common challenges in existing code search datasets:
  1. C1
    Unrealistic Queries. Datasets like CodeSearchNet (Husain et al., 2019) contain general-purpose queries from GitHub, but these queries are either documentation strings or comments rather than natural language questions, limiting its practicality in real scenarios.
  2. C2
    Mismatched Codes. CoSQA contains realistic queries from Bing search engine. However, we found that there are around 51% queries paired with mismatching code.
  3. C3
    Single Code Pair. Existing code search datasets typically pair each query with a single code snippet. This one-to-one match limits the evaluation of queries that could correctly match multiple code snippets.
@@
To address these problems, we develop a new benchmark CoSQA+, pairing high-quality queries with multiple high-quality codes. The queries come from CoSQA, and the codes include snippets, blocks and functions from the filtered StaQC and CSN datasets. The candidate pairs are formed by pairing the queries with the codes through multiple models. Utilizing powerful LLM, Claude 3 Sonnet, we label the candidate pairs automatically with CoT prompts. Given the limited size of the code pool, some queries fail to be matched with suitable code snippets. We use GPT-4o to generate exact matching codes for these queries and manually verify their accuracy. Furthermore, a new metric, MMRR, is proposed to evaluate models’ performance on one-to-N code search.
@@
We conduct four experiments to assess the quality of CoSQA+. (1) In a quality comparison between CoSQA+ and CoSQA, for the 1000 queries, 62.9% of the paired codes from CoSQA+ are selected as better. (2) When CodeBERT is fine-tuned on CoSQA+, it demonstrates superior performance in the CSN Python than when fine-tuned on CoSQA, with the MMRR of 0.902 for CoSQA+ versus 0.850 for CoSQA. (3) Automated Claude 3 Sonnet annotation yields performance close to human levels, with a Krippendorff’s Alpha of 0.628 and an accuracy of 84% in exact match conditions. (4) The MMRR metric proves to be highly reliable and stable for evaluating the effectiveness of multi-choice code search on CoSQA+, as evidenced by Cronbach’s Alpha of 0.956 and variance of 0.0004.
@@
2 Related Work
2.1 Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search. Husain et al. Husain et al. (2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA (Huang et al., 2021) and CS1QA (Lee et al., 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al., 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality. More recently, Li et al. (Li et al., 2024) introduce ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This SOTA benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
2.2 Code Search Methods
The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models. IR based methods extract keywords from queries and search for these keywords in code base to retrieve the most relevant code snippets (Bajracharya et al., 2006; Di Grazia and Pradel, 2023; Zhang et al., 2010). However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly. Fortunately, DL based methods receive great attention in NLP and achieve significant breakthroughs in code search research. The CodeSearchNet (Husain et al., 2019) dataset paves the way for CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), both of which are bimodal pre-trained BERT (Devlin et al., 2018) based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) (Clark et al., 2020) objective. UniXcoder (Guo et al., 2022) unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment. Recent T5 (Raffel et al., 2020) based models like CodeT5 (Wang et al., 2021) and CodeT5+ (Wang et al., 2023) show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets.
As large-scale models gain prominence, GPT (Radford et al., 2018) based models such as Codex (Chen et al., 2021), Claude 3 (Anthropic, 2024) and GPT-4 (Achiam et al., 2023) are also employed in the field of code search, and exhibit remarkable performance.
@@ 3 CoSQA+
3.1 Query and Code Collection
We take CoSQA as the query source, while CodeSearchNet Corpus (Husain et al., 2019) and StaQC (Yao et al., 2018) serve as the code sources. This section details the code data collection process, including the selection of appropriate code formats, sources, and a cleaning tool to promote the quality of the dataset.
@@ 3.1 Query and Code Collection
The selection of code format is important for constructing a query-code matching dataset (Huang et al., 2021). Different formats include a statement (Yin et al., 2018), a code snippet or block (Yao et al., 2018), and a function (Husain et al., 2019). CoSQA simplifies the task and adopts a complete Python function as the answer, while we also adopt a code snippet or block for the following reasons. (1) Adopting a more inclusive code format allows for the collection of a broader dataset, which in turn can create a more comprehensive dataset. (2) This approach facilitates the evaluation of a model’s search capabilities across different code formats. (3) Diverse code examples play a crucial role in enhancing search accuracy by showcasing the variety of coding styles, syntax, and structural semantics  (Oliveira et al., 2023).
@@
To the best of our knowledge, CodeSearchNet and StaQC are the only origin datasets specifically designed for Python in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages (Husain et al., 2019). In CoSQA+, we keep all 1,156,085 Python functions. In addition, around 148K Python code snippets from StaQC are added to the candidate code pool of CoSQA+.
@@
The 148K Python code base in StaQC contains many non-Python code snippets. Guesslang [^1] https://pypi.org/project/guesslang/ is a programming language detection tool based on Tensorflow. We utilize Guesslang to identify and filter out non-Python codes. The filtered-out codes can be categorized into four types: (1) incomplete snippets, (2) irrelevant fields, (3) codes written in other programming languages, and (4) others. More explanation and examples are shown in Appendix A. After cleaning, we retain 79k Python codes in StaQC.
3.2 Candidate Query-Code Pairs Construction
@@
High similarity between query and code is fundamental for a high-quality dataset (Sun et al., 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022) and CodeT5+ 110M embedding (Wang et al., 2023). By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
@@ 3.3 Model Annotation
The traditional data annotation process heavily relies on manual labor, demanding a deep understanding of Python and its libraries. Even experienced programmers need to use search tools to annotate query-code pairs, which is inefficient and time-consuming. In contrast, advanced language models like Claude 3 Sonnet (Anthropic, 2024) offer an alternative. With comprehensive training and robust computational abilities, Claude 3 Sonnet can promote quality and accelerate the annotation process. Recognizing this potential, we implement a Chain of Thought (CoT) method to automate the annotation of each query-code pair. The CoT method integrates the specific guidelines directly into the model’s prompting mechanism. The performance of LLMs annotation is further evaluated in the experiment section.
3.3.1 Annotation Guidelines
@@
We compare our annotation guidelines with other datasets, such as CodeSearchNet (Husain et al., 2019) and CoSQA (Huang et al., 2021), and identify the key difference: these datasets deem codes correct if they meet just over half of the query’s demands. In contrast, our dataset adopts stricter labeling criteria, where only codes that fully satisfy the query’s requirements are labeled as correct. These stricter standards are implemented to better quantify the extent to which code meets the requirements of a query, thereby reducing inconsistencies in query-code pair annotations. Furthermore, stricter standards better align with the programmers’ requirements for retrieving more accurately matched query-code pairs in real-world scenarios.
3.3.2 Annotation
@@
3.4 Missing Code Generation
@@ 4 Evaluation
4.1 RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior performance compared to CoSQA.
@@ 4.1 RQ1: Quality Comparison between CoSQA+ and CoSQA
In the first experiment, we randomly select 1000 query-code pairs labeled “1” from both CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, from which we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced human programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, we fine-tune CodeBERT with CoSQA+ and CoSQA separately, testing the models on the CSN Python with MRR and MMRR as evaluation metrics. For CoSQA+, the dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division. From the CSN benchmark, a rigorously human-verified code search benchmark, we extract Python language test samples, resulting in 399 query-code pairs for 95 queries, referred to as CSN Python.
Fine-tuned Data MRR MMRR
CoSQA 0.939 0.850
CoSQA+ 0.962 0.902
Table 2: Performance of CodeBERT fine-tuned with different data. The test set is CSN Python.
@@
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 62.9% of the better codes are from CoSQA+, while 37.1% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, CodeBERT fine-tuned with CoSQA+ outperforms that fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in MMRR is more significant than the improvement in MRR.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
4.2 RQ2: Annotation Performance of LLMs
@@
This experiment can be divided into two parts. In the first part, we utilize Krippendorff’s alpha coefficients (Krippendorff, 1980) $\alpha$ to measure the degree of agreement among LLM annotators and human annotators. For the second part, we use the metric Accuracy to evaluate the performance of humans and different LLMs.
@@
For the models evaluation on Krippendorff’s alpha coefficients (Krippendorff, 1980) $\alpha$ and accuracy, we select six LLMs: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Claude 3 Sonnet, Claude 3 Opus and Llama3-70B-instruct, as shown in Table 3 and Table 4. These models label the same pairs under identical guidelines.
@@
The human annotator’s Krippendorff’s Alpha Coefficients are 0.354 and 0.607 for 50% match and exact match, respectively. As shown in Table 4, Claude 3 Sonnet outperforms the other models in this metric and is close to human levels. Therefore, it is selected to perform the three rounds of labeling the query-code pairs for the data annotation part. To guarantee the reliability and stability of labeling, Krippendorff’s alpha coefficients (Krippendorff, 1980) are also calculated between different labeling rounds. Claude 3 Sonnet achieves a high score of 0.965 and 0.975, respectively.
@@
4.3 RQ3: Existing Methods Performance on CoSQA+
@@
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark. The primary metric for evaluation is the MMRR, which is specifically designed for tasks where multiple correct solutions may exist for a given query. The MMRR better reflects the complexity and variability in code-related searches than the MRR does. The tested methods encompass both traditional information retrieval techniques like Lucene, Bag of Words(BoW) (Zhang et al., 2010) and advanced models like CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022), CodeT5+ 110M embedding (Wang et al., 2023) and text-embedding-3-large [^2] https://openai.com/index/new-embedding-models-and-api-updates/.
@@
Advanced models like CodeBERT, UniXcoder, CodeT5+ and text-embedding-3-large significantly improve on MMRR and MRR, owing to the training on a mix of natural and program languages. Text-embedding-3-large, as a text embedding model rather than one specifically designed for code search, achieves the best results among all the methods. This may be due to its higher dimensions and its superior embedding quality.
@@
4.4 RQ4: Is MMRR a Stable and Reliable Metric in Multi-Choice Code Search?
@@
We employ Cronbach’s Alpha Tavakol and Dennick (2011) to measure the reliability of CoSQA+ in the MMRR. Additionally, we use average variance to evaluate the stability of the MMRR across different test sets and models. For stability evaluation, we calculate the average variance Fisher (1936) of the MMRR across different models and test sets. The variance is computed by each model performance across all test sets, and then these variances are averaged to provide a single stability measure.
@@
To evaluate the reliability and stability of MMRR in assessing models’ capability in multi-choice code search tasks, we exclude queries corresponding to only one correct code. The queries are divided into four distinct query bases, corresponding to the number of exactly matched codes (2, 3, 4, 5) for each query. The code base remains consistent with the original CoSQA+. Four different query bases form four test datasets with identical code base. We assess the multi-choice code search capabilities of three different models across these datasets.

### v2

@@ Abstract
Introduction
@@
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al. 2013; Sim et al. 2011; Stolee, Elbaum, and Dobos 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al. 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. For the practical usability of neural code search, real user queries are crucial to a high-quality dataset  (Sun et al. 2022). The key challenge is how to pair real-world queries with codes for a high-quality code search benchmark.
However, there are some common challenges in existing code search datasets. Datasets like CodeSearchNet (Husain et al. 2019) contain general-purpose queries from GitHub, but these queries are either documentation strings or comments rather than natural language questions. Models trained on these unrealistic queries often fail to generalize to real-world use cases, leading to poor performance in practical applications (Liu et al. 2009; Xie et al. 2023). CoSQA contains realistic queries from the Bing search engine, but we found that around 51% of queries are paired with mismatching code, which adversely affects the training of code search models. When code snippets do not correspond accurately to the query descriptions, this leads to excessive noise in the dataset, complicating the model’s ability to retrieve relevant code (Li, Zhou, and Shen 2024; Sun et al. 2022). Additionally, existing code search datasets typically pair each query with a single code snippet, which limits the evaluation of queries that could match multiple code snippets. Finally, building query-code datasets involves laboriously matching and labeling real search queries with accurate code. Despite employing 100 human annotators to cover 20,604 queries, CoSQA only managed to generate 6,267 unique code snippets, with half of the queries failing to match the correct code.
@@
To address these problems, we develop a new benchmark CoSQA+, pairing high-quality queries with multiple high-quality codes. The queries come from CoSQA, and the codes include snippets, blocks and functions from the filtered StaQC and CodeSearchNet datasets. The candidate pairs are formed by pairing the queries with the codes through multiple models. Utilizing powerful Claude 3 Sonnet, we automatically label the candidate pairs with CoT prompts. Given the limited size of our code pool, some queries fail to be matched with suitable code snippets. We use GPT-4o to generate exact matching codes for these queries and manually verify their accuracy. Furthermore, a new metric, MMRR, is proposed to evaluate models’ performance on one-to-N code search.
@@
We conduct four experiments to assess the quality of CoSQA+. (1) In a quality comparison between CoSQA+ and CoSQA, for the 1000 queries, 62.9% of the paired codes from CoSQA+ are selected as better. (2) When CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings are fine-tuned on CoSQA+, they all perform better on the CSN Python dataset compared to being fine-tuned on CoSQA. (3) Automated Claude 3 Sonnet annotation yields performance close to human levels, with a Krippendorff’s Alpha of 0.628 and an accuracy of 84% in exact match conditions. (4) The MMRR metric proves to be highly reliable and stable for evaluating the effectiveness of multi-choice code search on CoSQA+, as evidenced by Cronbach’s Alpha of 0.956 and variance of 0.0004.
@@
Related Work
Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search. Husain et al. (Husain et al. 2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA (Huang et al. 2021) and CS1QA (Lee, Seonwoo, and Oh 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al. 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality. More recently, Li et al. (Li et al. 2024) introduce ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This SOTA benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
Code Search Methods
The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models. IR based methods extract keywords from queries and search for these keywords in code base to retrieve the most relevant code snippets (Bajracharya et al. 2006; Di Grazia and Pradel 2023; Zhang, Jin, and Zhou 2010). However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly. Fortunately, DL based methods receive great attention in NLP and achieve significant breakthroughs in code search research. The CodeSearchNet (Husain et al. 2019) dataset paves the way for CodeBERT (Feng et al. 2020) and GraphCodeBERT (Guo et al. 2020), both of which are bimodal pre-trained BERT (Devlin et al. 2018) based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) (Clark et al. 2020) objective. UniXcoder (Guo et al. 2022) unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment. Recent T5 (Raffel et al. 2020) based models like CodeT5 (Wang et al. 2021) and CodeT5+ (Wang et al. 2023) show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets. While the two models achieve high performances in the CSN benchmark, GraphCodeBERT (Guo et al. 2020) demonstrates superior results by harnessing data flow encoded in the Abstract Syntax Tree (AST) of code to enrich the structural information during pre-training.
As large-scale models gain prominence, GPT (Radford et al. 2018) based models such as Codex (Chen et al. 2021), Claude 3 (Anthropic 2024) and GPT-4 (Achiam et al. 2023) are also employed in the field of code search, and exhibit remarkable performance.
@@
Query and Code Collection
We take CoSQA as the query source, while CodeSearchNet Corpus (Husain et al. 2019) and StaQC (Yao et al. 2018) serve as the code sources. This section details the code data collection process, including the selection of appropriate code formats, sources, and a cleaning tool to promote the quality of the dataset.
Query Source
CoSQA queries are based on real user queries collected from Microsoft’s Bing search engine logs. We use these queries as the source for the following reasons. (1) We concentrate on code search in Python, and the filtered queries specifically target Python code search intent (Huang et al. 2021). (2) Utilizing real-world queries allows the code search model to be trained on data that reflects actual developer behavior. This helps the model effectively handle the diverse ways developers express their coding inquiries in real-world scenarios. (3) By using realistic queries, the code search model better meets user expectations, enhancing overall satisfaction. Developers are more likely to find relevant results when the queries they input closely align with their thought processes and communication about code (Li et al. 2020).
@@
The selection of code format is important for constructing a query-code matching dataset (Huang et al. 2021). Different formats include a statement (Yin et al. 2018), a code snippet or block (Yao et al. 2018), and a function (Husain et al. 2019). CoSQA simplifies the task and adopts a complete Python function as the answer, while we also adopt a code snippet or block for the following reasons. (1) Adopting a more inclusive code format allows for the collection of a broader dataset, which in turn can create a more comprehensive dataset. (2) This approach facilitates the evaluation of a model’s search capabilities across different code formats. (3) Diverse code examples play a crucial role in enhancing search accuracy by showcasing the variety of coding styles, syntax, and structural semantics  (Oliveira et al. 2023).
@@
To the best of our knowledge, CodeSearchNet and StaQC are the only origin datasets specifically designed for Python in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages (Husain et al. 2019). In CoSQA+, we keep all 1,156,085 Python functions. In addition, around 148K Python code snippets from StaQC are added to the candidate code pool of CoSQA+.
@@
The 148K Python code base in StaQC contains many non-Python code snippets. Guesslang [^2] https://pypi.org/project/guesslang/ is a programming language detection tool based on Tensorflow. We utilize Guesslang to identify and filter out non-Python codes. The filtered-out codes can be categorized into four types: (1) incomplete snippets, (2) irrelevant fields, (3) codes written in other programming languages, and (4) others. More explanation and examples are shown in Appendix Figure 1. After cleaning, we retain 79k Python codes in StaQC.
Candidate Query-Code Pairs Construction
@@
High similarity between query and code is fundamental for a high-quality dataset (Sun et al. 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT (Feng et al. 2020), UniXcoder (Guo et al. 2022) and CodeT5+ 110M embedding (Wang et al. 2023). By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
@@
The traditional data annotation process heavily relies on manual labor, demanding a deep understanding of Python and its libraries. Even experienced programmers need to use search tools to annotate query-code pairs, which is inefficient and time-consuming. In contrast, advanced language models like Claude 3 Sonnet (Anthropic 2024) offer an alternative. With comprehensive training and robust computational abilities, Claude 3 Sonnet can promote quality and accelerate the annotation process. Recognizing this potential, we implement a Chain of Thought (CoT) method to automate the annotation of each query-code pair. The CoT method integrates the specific guidelines directly into the model’s prompting mechanism. The performance of LLMs annotation is further evaluated in the experiment section.
Annotation Guidelines
@@
We compare our annotation guidelines with other datasets, such as CodeSearchNet (Husain et al. 2019) and CoSQA (Huang et al. 2021), and identify the key difference: these datasets deem codes correct if they meet just over half of the query’s demands. In contrast, our dataset adopts stricter labeling criteria, where only codes that fully satisfy the query’s requirements are labeled as correct. These stricter standards are implemented to better quantify the extent to which code meets the requirements of a query, thereby reducing inconsistencies in query-code pair annotations. Furthermore, stricter standards better align with the programmers’ requirements for retrieving more accurately matched query-code pairs in real-world scenarios.
Annotation
@@
Missing Code Generation
@@
RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
@@
In the first experiment, we randomly select 1000 query-code pairs labeled “1” from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, we fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with MRR and MMRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
Model Metric Fine-tune Dataset
None CoSQA CoSQA+
CodeBERT MRR 0.952 0.939 0.961
MMRR 0.882 0.844 0.899
UniXcoder MRR 0.891 0.951 0.954
MMRR 0.711 0.857 0.884
CodeT5+ MRR 0.794 0.938 0.953
embedding MMRR 0.504 0.860 0.889
Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
@@
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 62.9% of the better codes are from CoSQA+, while 37.1% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in MMRR is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
RQ2: Annotation Performance of LLMs
@@
This experiment can be divided into two parts. In the first part, we utilize Krippendorff’s alpha coefficients (Krippendorff 1980) $\alpha$ to measure the degree of agreement among LLM annotators and human annotators. For the second part, we use the metric Accuracy to evaluate the performance of humans and different LLMs.
@@
For the models evaluation on Krippendorff’s alpha coefficients (Krippendorff 1980) $\alpha$ and accuracy, we select six LLMs: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Claude 3 Sonnet, Claude 3 Opus and Llama3-70B-instruct, as shown in Table 3 and Table 4. These models label the same pairs under identical guidelines.
@@
The human annotator’s Krippendorff’s Alpha Coefficients are 0.354 and 0.607 for 50% match and exact match, respectively. As shown in Table 4, Claude 3 Sonnet outperforms the other models in this metric and is close to human levels. Therefore, it is selected to perform the three rounds of labeling the query-code pairs for the data annotation part. To guarantee the reliability and stability of labeling, Krippendorff’s alpha coefficients (Krippendorff 1980) are also calculated between different labeling rounds. Claude 3 Sonnet achieves a high score of 0.965 and 0.975, respectively. Other models show little difference in scores under the 50% match criteria, while varying significantly under the exact match criteria. This is because of the variations in their understanding of how codes can fully match the query.
@@
RQ3: Existing Methods Performance on CoSQA+
@@
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark. The primary metric for evaluation is the MMRR, which is specifically designed for tasks where multiple correct solutions may exist for a given query. The MMRR better reflects the complexity and variability in code-related searches than the MRR does. The tested methods encompass both traditional information retrieval techniques like Lucene, Bag of Words(BoW) (Zhang, Jin, and Zhou 2010) and advanced models like CodeBERT (Feng et al. 2020), UniXcoder (Guo et al. 2022), CodeT5+ 110M embedding (Wang et al. 2023) and text-embedding-3-large [^3] https://openai.com/index/new-embedding-models-and-api-updates/.
@@
Advanced models like CodeBERT, UniXcoder, CodeT5+ and text-embedding-3-large significantly improve on MMRR and MRR, owing to training on a mix of natural and program languages. Text-embedding-3-large, as a text embedding model rather than one specifically designed for code search, achieves the best results among all the methods. This may be due to its higher dimensions and superior embedding quality. CodeBERT achieves results second only to Text-embedding-3-large. This performance is probably due to its encoder-only architecture and extensive pre-training on large pairs of natural and programming languages, especially Python, making it highly effective for code understanding.
@@
RQ4: Is MMRR a Stable and Reliable Metric in Multi-Choice Code Search?
@@
We employ Cronbach’s Alpha (Tavakol and Dennick 2011) to measure the reliability of CoSQA+ in the MMRR. Additionally, we use average variance to evaluate the internal consistency and stability of the MMRR across different test sets and models. For stability evaluation, we calculate the average variance (Fisher 1936) of the MMRR across different models and test sets. The variance is computed by each model performance across all test sets, and then these variances are averaged to provide a single stability measure.
@@
To evaluate the reliability and stability of MMRR in assessing models’ capability in multi-choice code search tasks, we exclude queries corresponding to only one correct code. The queries are divided into four distinct query bases, corresponding to the number of exactly matched codes (2, 3, 4, 5) for each query. The code base remains consistent with the original CoSQA+. This stratification allows us to explore how MMRR performs across queries of varying complexity and choice. Four different query bases form four test datasets with identical code base. We assess the multi-choice code search capabilities of three different models across these datasets.